In a survey on downloading​ music, a poll asked 728 internet users if they​ "ever downloaded music from an Internet site that was not authorized by a record​ company, or​ not," and 16​% responded​ "yes." Construct a 99

​% confidence interval for the true proportion of internet users who have downloaded music from an Internet site that was not authorized. What is the confidence​ interval? Select the correct choice below​ and, if​ necessary, fill in the answer boxes within your choice.
Answer #1

solution Here n=728, P = 16% = 0.16 At 99% confidence, za = 2-576 Hargin of erra (4 Ð) = Zoy ( [8.16 (1-0.16 = 2-576 728 = 0.